what is the slope and y-intercept for 4-1/3y=5x
Answer by nerdybill(7384)

Put:
4-1/3y=5x
int "slope-intercept" form:
y = mx + b
where
m is the slope
b is the y-intercept
.
4-1/3y=5x
4=1/3y+5x
-5x+4=1/3y
-15x+12=y
.
slope is -15
y-intecept is at (0,12)